OpenStudy (anonymous):

Clarification: If f(x)= 6x^3=9x^2-x +C and f(-3)=1; C=-245?

OpenStudy (anonymous):

I think you have too many equals signs

OpenStudy (anonymous):

This is how it is on the paper. I solved it with: f(x) = 6x^3 + 9x^2-X+C 162+81- (-3) +C F(-3) =1= 246+C -246+1=-245 -245=C

OpenStudy (anonymous):

6(-3)^3 = -162

OpenStudy (anonymous):

??? I don't understand

OpenStudy (anonymous):

-162+81+3+c=1 78+c=1 c=-77

OpenStudy (anonymous):

That's not an answer choice. I have C=-245; c=79; c=31; and c=-83

OpenStudy (anonymous):

sorry -78+c=1 c=79

OpenStudy (anonymous):

What did I do wrong? May you explain?

OpenStudy (anonymous):

when you cubed the -3, so changed it to positive instead of leaving it as negative

OpenStudy (anonymous):

Okay thank you... Can you help me on one more thing please?

OpenStudy (anonymous):

I found the Y-intercepts. However for the x-intercepts of f(x)= (x+1) (x-6) (x-1)^2 I'm not sure how to find the x- intercepts.

OpenStudy (anonymous):

the x-intercepts are when y=0 (x+1)(x-6)(x-1)(x-1)=0 the zero product rule says if any or all are 0, it is true so x={-1,1,6}

OpenStudy (anonymous):

Okay so i solved it like this: 0= -1(x+1) (x-6) (x-1)^2

OpenStudy (anonymous):

why 0=-1(x+1)(x-6)(x-1)^2 why the -1?

OpenStudy (anonymous):

Well I did some research and someone did it that way I ended up getting the same answer as you. What did you do?

OpenStudy (anonymous):

you did it fine, just when you did 6(-3)^3 you got a positive answer instead of a negative answer. (-3)^3 is -27 not +27

OpenStudy (anonymous):

I'm talking about for the intercepts?

OpenStudy (anonymous):

the -1 doesn't play any role in finding the intercepts so putting it there wont change the answer but it shouldn't be there

OpenStudy (anonymous):

So how would you have solved it?

OpenStudy (anonymous):

the x-intercepts are when y=0 (x+1)(x-6)(x-1)(x-1)=0 the zero product rule says if any or all are 0, it is true so x={-1,1,6} I did it the same way as you but you put -1(x+1)(x-6)(x-1)(x-1)=0 [ the -1 in front] which doesn't change the answer but shouldn't be there

OpenStudy (anonymous):

Okay sorry for all the trouble. Thanks.

OpenStudy (anonymous):

Why does -6 turn to 6 then?

OpenStudy (anonymous):

(x-6)=0 so x=6 (x+1)=0 so x=-1 (x-1)=0 so x=1

OpenStudy (anonymous):

Right! I feel so dumb. Thank you!